how do you find the ordered pair for the equation y=-9x+6?
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
y(x) = -9x + 6
---
you can find an infinite number of ordered pairs by substitution:
y(1) = -9*1 + 6 = -3 : (1, -3)
y(2) = -9*2 + 6 = -12 : (2, -12)
and so on
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

how do you find the ordered pair for the equation y=-9x+6?
----------
x is the independent variable: its value is your choice.
----
Example: let x = 5
----
Now, solve for "y" :: y = -9*5 + 6 = -45+6 = -39
====================
Cheers,
Stan H.
====================